### v2

@@ Discussion
In the present manuscript, we have presented the capabilities of explainable artificial intelligence (XAI) to estimate the importance of each grid point in a turbulent channel for the prediction of its future states. Then, new coherent structures are defined based on these importance scores, namely the SHAP values. The SHAP-based structures have been analyzed from a physical point of view by comparing them against other coherent structures widely studied in the turbulence literature: intense Reynolds-stress structures (Q events) [35], streaks [18] and vortices [20]. Near the wall, a strong correlation between the SHAP structures and the sweeps can be observed, especially in the regions where the sweeps collide with high-velocity streaks. Then, for a wall-normal distance $y^{+}\approx 15$, the streaks increase their presence in the flow, agreeing with the SHAP structures. Finally, farther from the wall, the presence of the streaks is lower and the SHAP structures mainly correlate with the ejections. Nevertheless, although the vortices increase their agreement with the SHAP structures near the mid-plane of the channel, their importance is still lower than that of the Q events.
@@
For the past 140 years, the wall-bounded-turbulence analysis has focused on the assessment of partial results in certain regions of the flow. The SHAP structures are the first method capable of addressing this question globally and objectively, identifying the regions of highest importance without any preconceived bias. Although the SHAP structures partly correlate with the classically studied ones in different regions of the channel, this agreement is not perfect. Since the SHAP structures objectively identify the most important regions of the flow, analyzing the differences between SHAP and other structures can provide an excellent venue to deepen our insight into wall-bounded turbulence. It is also important to highlight that in this work the importance is defined based on a surrogate model which predicts a state of the flow 5 viscous times into the future. Other possible models can be developed, e.g. for predictions with longer time horizons (characteristic of the outer region) or of the wall-shear stress (where different SHAP structures may be identified). The different questions that can be addressed with various models will be addressed in future studies.
@@ Methods
The velocity fields of turbulent channel flow used in this study were calculated by numerically integrating the Navier–Stokes equations using the LISO code. This direct-numerical-simulation (DNS) code has successfully been used for solving some of the largest simulations in the field of wall-bounded turbulence [36, 15]. The integration strategy of LISO is based on the spectral method proposed by Kim et al. [37] using a seven-point compact-finite-difference scheme in the wall-normal direction with fourth-order consistency and extended spectral-like resolution [38]. Concerning the temporal discretization, a third-order semi-implicit Runge–Kutta scheme is employed [39]. Constant grid spacings of approximately $8.2$ and $4.1$ viscous units are used in the streamwise and spanwise directions, respectively. The wall-normal grid spacing is adjusted along the wall-normal distance to maintain the ratio between the grid spacing and the local isotropic Kolmogorov scale ($\eta$) constant: $\Delta y=1.5\eta$. The local isotropic Kolmogorov scale is defined as $\eta=\left(\nu^{3}/\varepsilon\right)^{1/4}$, where $\varepsilon=2\nu\left(\partial u_{i}/\partial x_{j}\times\partial u_{j}/\partial x_{i}\right)$ is the viscous dissipation rate and $\nu$ the kinematic viscosity.
A turbulent channel with dimensions $8\pi h\times 2h\times 3\pi h$ in the streamwise, wall-normal and spanwise dimensions is simulated, where $h$ is half the distance between the channel walls. These dimensions are selected to ensure that the domain is large enough to include multiple coherent structures of each type in both the streamwise and spanwise directions, as can be observed in Figure 2. The previous grid spacing and channel size generate a mesh containing $384\times 201\times 288$ points in the streamwise, wall-normal and spanwise directions, which matches the initial and final sizes of the data used in the U-net presented in Figure 1.
@@
Regarding the deep-learning model, a U-net architecture [33], similar to the one used in Cremades et al. [27], is selected for predicting the future state of the flow. The U-net takes the three-dimensional fields of the 3 velocity fluctuations at an instant $t$ as input and predicts the three fluctuation fields five viscous times into the future (the viscous time is defined in terms of $\nu$ and $u_{\tau}^{2}$). Note that, in a previous work [27], we show that these results were consistent for a range between one to five viscous times. A schematic representation of the U-net can be observed in the central block of Figure 1. The architecture comprises four layers, where the first layer has a size of $201\times 288\times 384$ points, which is padded into the shape $201\times 318\times 414$ (blue arrow) in the first operation, exploiting the periodicity of the channel to avoid any possible error in the edges of the convolution, and cropped back (light blue arrows) in the last operation. In the encoder, two three-dimensional convolutions are applied with $32$ filters each one; in the decoder, the first three-dimensional convolution with $32$ filters is followed by a convolution with 3 filters for the output field adapting the information to the output field. The second, third, and fourth layers exhibit field sizes of $100\times 159\times 207$, $50\times 79\times 103$, and $25\times 39\times 51$, with $64$, $128$ and $256$ filters respectively. The different levels contain two three-dimensional convolutions (violet arrows) in the encoder and one in the decoder. In the encoder, the information is compressed from one level to the following through a three-dimensional mean pooling (pink arrows), and it is expanded in the decoder using transposed three-dimensional convolutions (gray arrows). In addition, the equivalent level in the encoder and the decoder are connected with concatenations that connect the final field of the encoder level with the output of the transposed convolutions, doubling the number of channels of the tensors. The U-net architectures are commonly used for image segmentation as they exploit the spatial patterns of the input field and they keep the low- and high-frequency information of the initial field [40, 41, 42].
@@
The U-net is trained on a database comprising a total of 10000 instantaneous flow fields of a turbulent channel at a friction Reynolds number of $\rm{Re}_{\tau}=125$, using 80% of them for training and 20% for validation until the relative error between the predicted output and the ground truth is approximately 1% in the three velocity components (1.15% in the streamwise, 0.98% in the wall-normal, and 1.08% in the spanwise velocity fluctuations). Then a total of 8000 fields are used for calculating the SHAP values of the flow. For the training process, a RMSprop gradient-descent algorithm was used [43], with a learning rate of $5\times 10^{-5}$ and a momentum of $0.9$.
@@
where $\phi_{0}$ is the MSE of the prediction of the reference (non-informative) field, $\phi_{\rm{ji}}$ is the SHAP value relative to the velocity component $j$ and the grid point $i$ and $z_{ji}$ is a boolean value, 00 if the grid point value is taken from the reference field (non-informative) or $1$ if it is taken from the input field.
@@
The minimum representative data used in this study will be made available open access at: https://github.com/KTH-FlowAI. For the complete database, please contact the authors.
@@
The codes used for this work will be made available open access at: https://github.com/KTH-FlowAI.
@@ Acknowledgments
The deep-learning-model training was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Berzelius (NSC), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. Part of the postprocess was made in the supercomputer Sirius of the Universitat Politècnica de València. SH has the support of grant PID2021-128676OB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”, by the European Union. RV acknowledges the financial support from ERC grant no. 2021-CoG-101043998, DEEPCONTROL. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@
Cremades, A.: Conceptualization, Methodology, Software, Validation, Investigation, Writing - Original Draft, Visualization. Hoyas, S.: Conceptualization, Data curation, resources, Writing - Original Draft, Funding acquisition. Vinuesa, R.: Initial idea, conceptualization, project definition, methodology, resources, Writing - Original Draft, Supervision, Project administration, Funding acquisition.

### v3

@@ Discussion
In the present manuscript, we have presented the capabilities of explainable artificial intelligence (XAI) to estimate the importance of each grid point in a turbulent channel up to a friction Reynolds number ${\rm Re}_{\tau}=550$ for the prediction of its future states. Then, new coherent structures are defined based on these importance scores, namely the SHAP values. The SHAP-based structures have been analyzed from a physical point of view by comparing them against other coherent structures widely studied in the turbulence literature: intense Reynolds-stress structures (Q events) [36], streaks [17] and vortices [19]. Near the wall, a strong correlation between the SHAP structures and the sweeps can be observed, especially in the regions where the sweeps collide with high-velocity streaks. Then, for a wall-normal distance $y^{+}\approx 15$, the streaks increase their presence in the flow, agreeing with the SHAP structures. Finally, farther from the wall, the presence of the streaks is lower and the SHAP structures mainly correlate with the ejections, although for ${\rm Re}_{\tau}=550$ the separation of scales is higher and sweeps are also present. Nevertheless, although the vortices increase their agreement with the SHAP structures near the mid-plane of the channel, their importance is still lower than that of the Q events.
@@
For the past 140 years, the wall-bounded-turbulence analysis has focused on the assessment of partial results in certain regions of the flow. The SHAP structures are the first method capable of addressing this question globally and objectively, identifying the regions of highest importance without any preconceived bias. The SHAP values are used to identify these high-importance regions by analyzing the most sensitive input features for a mathematically defined question: which regions minimize the error of the model predictions? Although the SHAP structures partly correlate with the classically studied ones in different regions of the channel, this agreement is not perfect. Since the SHAP structures objectively identify the most important regions of the flow, analyzing the differences between SHAP and other structures can provide an excellent venue to deepen our insight into wall-bounded turbulence, uncovering the most important causal relations in the system. An example of application of the present SHAP framework to another flow case, namely the flow around a wall-mounted square obstacle [23, 38], can be found in the Supplementary Material. In this case, the SHAP values highlight the influence of the wall and the obstacle, reducing the importance of the streaks , identifying a similar importance of the Q events and a low importance of the vortices.. The definition of causally important regions is a change of paradigm in the analysis of physical problems, and can be replicated in many other fields such as thermal fields or cavitation. Furthermore, it can help to develop much more efficient control strategies [34]. In addition, the methodology has been proved to be scalable with the friction Reynolds number, producing results directly related to the wall-normal distance independently of the Reynolds number. It is also important to highlight that in this work the importance is defined based on a surrogate model which predicts a state of the flow 5 viscous times into the future. The Supplementary Material also presents results for a time interval $\Delta t^{+}=10$, showing high levels of agreement with the analysis for $\Delta t^{+}=5$. Other possible models can be developed, e.g. for predictions with much longer time horizons (characteristic of the outer region) or of the wall-shear stress (where different SHAP structures may be identified). In fact, the SHAP structures for the evolution of the vorticity are presented in the Supplementary Material. The different questions that can be formulated with various models will be addressed in future studies.
@@ Methods
The velocity fields of turbulent channel flow used in this study were calculated by numerically integrating the Navier–Stokes equations using the LISO code. This direct-numerical-simulation (DNS) code has successfully been used for solving some of the largest simulations in the field of wall-bounded turbulence [14, 39]. The integration strategy of LISO is based on the spectral method proposed by Kim et al. [40] using a seven-point compact-finite-difference scheme in the wall-normal direction with fourth-order consistency and extended spectral-like resolution [41]. Concerning the temporal discretization, a third-order semi-implicit Runge–Kutta scheme is employed [42]. Constant grid spacings of approximately $8.2$ and $4.1$ viscous units are used in the streamwise and spanwise directions, respectively for the ${\rm Re}_{\tau}=125$ case, and $8.9$ and $4.3$ for the ${\rm Re}_{\tau}=550$ case. The wall-normal grid spacing is adjusted along the wall-normal distance to maintain the ratio between the grid spacing and the local isotropic Kolmogorov scale ($\eta$) constant: $\Delta y=1.5\eta$. The local isotropic Kolmogorov scale is defined as $\eta=\left(\nu^{3}/\varepsilon\right)^{1/4}$, where $\varepsilon=2\nu\left(\partial u_{i}/\partial x_{j}\cdot\partial u_{j}/\partial x_{i}\right)$ is the viscous dissipation rate and $\nu$ the kinematic viscosity.
A turbulent channel with dimensions $8\pi h\times 2h\times 3\pi h$ in the streamwise, wall-normal and spanwise dimensions is simulated for ${\rm Re}_{\tau}=125$ and $2\pi h\times 2h\times\pi h$ for ${\rm Re}_{\tau}=550$, where $h$ is half the distance between the channel walls. These dimensions are selected to ensure that the domain is large enough to include multiple coherent structures of each type in both the streamwise and spanwise directions, as can be observed in Figure 2. The previous grid spacing and channel size generate a mesh containing $384\times 201\times 288$ points in the streamwise, wall-normal and spanwise directions for ${\rm Re}_{\tau}=125$ and $384\times 251\times 384$ for ${\rm Re}_{\tau}=550$, which matches the initial and final sizes of the data used in the U-net presented in Figure 1.
@@
Regarding the deep-learning model, a U-net architecture [32], similar to the one used in Cremades et al. [26], is selected for predicting the future state of the flow. The U-net takes the three-dimensional fields of the 3 velocity fluctuations at an instant $t$ as input and predicts the three fluctuation fields five viscous times into the future (the viscous time is defined in terms of $\nu$ and $u_{\tau}^{2}$). Note that, in a previous work [26], we showed that these results were consistent for a range between one to five viscous times. A schematic representation of the U-net can be observed in the central block of Figure 1. The architecture comprises four layers in the case of ${\rm Re}_{\tau}=125$, where the first layer has a size of $201\times 288\times 384$ points, which is padded into the shape $201\times 318\times 414$ (blue arrow) in the first operation, exploiting the periodicity of the channel to avoid any possible error in the edges of the convolution, and cropped back (light blue arrows) in the last operation. In the encoder, two three-dimensional convolutions are applied with $32$ filters each one; in the decoder, the first three-dimensional convolution with $32$ filters is followed by a convolution with 3 filters for the output field adapting the information to the output field. The second, third, and fourth layers exhibit field sizes of $100\times 159\times 207$, $50\times 79\times 103$, and $25\times 39\times 51$, with $64$, $128$ and $256$ filters respectively. The different levels contain two three-dimensional convolutions (violet arrows) in the encoder and one in the decoder. In the encoder, the information is compressed from one level to the following through a three-dimensional mean pooling (pink arrows), and it is expanded in the decoder using transposed three-dimensional convolutions (gray arrows). In addition, the equivalent level in the encoder and the decoder are connected with concatenations that connect the final field of the encoder level with the output of the transposed convolutions, doubling the number of channels of the tensors. For the ${\rm Re}_{\tau}=550$ case the number of filters is adjusted to $24$, $48$, $96$ and $192$ depending on the layer and the sizes of the tensors are modified according to the size of the input field. The U-net architectures are commonly used for image segmentation as they exploit the spatial patterns of the input field and they keep the low- and high-frequency information of the initial field [43, 44, 45].
@@
The U-net is trained on a database comprising a total of 10000 instantaneous flow fields of a turbulent channel at a friction Reynolds number of $\rm{Re}_{\tau}=125$ and $\rm{Re}_{\tau}=550$ respectively, using 80% of them for training and 20% for validation until the relative error between the predicted output and the ground truth is approximately 1% in the three velocity components (1.15% in the streamwise, 0.98% in the wall-normal, and 1.08% in the spanwise velocity fluctuations for $\rm{Re}_{\tau}=125$ and 0.74%, 1.15% and 0.88% for $\rm{Re}_{\tau}=550$). Then a total of 8000 fields are used for calculating the SHAP values of the flow. For the training process, a RMSprop gradient-descent algorithm was used [46], with a learning rate of $5\times 10^{-5}$ and a momentum of $0.9$.
@@
where $\phi_{0}$ is the MSE of the prediction of the reference (non-informative) field, $\phi_{\rm{ji}}$ is the SHAP value relative to the velocity component $j$ and the grid point $i$ and $z_{ji}$ is a boolean value, 0 if the grid point value is taken from the reference field (non-informative) or $1$ if it is taken from the input field.
@@
The minimum representative downsampled data used in this study will be made available open access at: https://github.com/KTH-FlowAI/XAI_turbulentchannel_3d_simplified.git. For the complete database, please contact the authors.
@@
The codes used for this work is available open access at: https://github.com/KTH-FlowAI/XAI_turbulentchannel_3d_simplified.git.
@@ Acknowledgments
The authors acknowledge Adrián Lozano-Durán and Álvaro Martínez-Sánchez for their support with the validation of the causal nature of the SHAP values. The deep-learning-model training was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Berzelius (NSC), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. Part of the postprocess was made in the supercomputer Sirius of the Universitat Politècnica de València. SH has the support of grant PID2021-128676OB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”, by the European Union. RV acknowledges the financial support from ERC grant no. 2021-CoG-101043998, DEEPCONTROL. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@
A. C.: Conceptualization, Methodology, Software, Validation, Investigation, Writing - Original Draft, Visualization. S. H.: Conceptualization, Data curation, resources, Writing - Original Draft, Funding acquisition. R. V.: Initial idea, conceptualization, project definition, methodology, resources, Writing - Original Draft, Supervision, Project administration, Funding acquisition.